# Notions of enriched purity

### v2

@@ 3. Enriched purity
Let $f\colon K\to L$ be a $\lambda$-filtered colimit in $\mathcal{K}^{\to}$ of some $(\lambda,\mathcal{E})$-pure morphisms $f_{i}\colon K_{i}\to L_{i}$, and consider any $g\colon A\to B$ in $\mathcal{K}_{\lambda}$. Since $g$ is a $\lambda$-presentable object of $\mathcal{K}^{\to}$, the map $\mathcal{K}(g,K)\colon\mathcal{K}(B,K)\to\mathcal{K}(A,K)$ is the $\lambda$-filtered colimit of the maps $\mathcal{K}(g,K_{i})$; the same applies for $L$ in place of $K$. Now, by $(\lambda,\mathcal{E})$-purity of each $f_{i}$, the $(\mathcal{E},\mathcal{M})$-factorization of each $\mathcal{K}(g,K_{i})$ is given by the object $\mathcal{P}(g,f_{i})$. We then obtain an induced $\lambda$-filtered diagram on the $\mathcal{P}(g,f_{i})$ whose colimit $X(g,f)$ induces the $(\mathcal{E},\mathcal{M})$-factorization of $\mathcal{K}(g,f)$; this is because both $\mathcal{E}$ and $\mathcal{M}$ are closed under $\lambda$-filtered colimits in $\mathcal{K}^{\to}$. But $\lambda$-filtered colimits commute with pullbacks in $\mathcal{V}$; thus $X(g,f)$ is the pullback of the colimit of $\mathcal{P}(g,L_{i})$ along $\mathcal{K}(A,f)$. Since this colimit coincides with $\mathcal{P}(g,L)$ then $X(g,f)\cong\mathcal{P}(g,f)$ and thus $f$ is $(\lambda,\mathcal{E})$-pure with respect to $g$. ∎
@@
Let $L\colon\mathcal{L}\to\mathcal{K}$ be the left adjoint to $F$; since $F$ preserves $\lambda$-filtered colimits then $L$ preserves the $\lambda$-presentable objects. Consider now a $(\lambda,\mathcal{E})$-pure morphisms $f\colon K\to L$ in $\mathcal{K}$ and any morphism $g\colon A\to B$ in $\mathcal{L}$ between $\lambda$-presentable objects. Then $\mathcal{L}(g,FL)\cong\mathcal{K}(Lg,L)$ and $\mathcal{L}(A,Ff)\cong\mathcal{K}(LA,f)$, so that
@@
as well as $r_{g,Ff}\cong r_{Lg,f}$. Now, since $A$ and $B$ are $\lambda$-presentable in $\mathcal{L}$, the morphism $Lg$ has $\lambda$-presentable domain and codomain in $\mathcal{K}$; thus we can apply the $(\lambda,\mathcal{E})$-purity of $f$ in $\mathcal{K}$ to deduce that $r_{Lg,f}$ lies in $\mathcal{E}$. Since that is isomorphic to $r_{g,Ff}$, the $(\lambda,\mathcal{E})$-purity of $Ff$ follows. ∎
@@
(1) It follows by the proposition above since $\mathcal{K}(C,-)$ preserves $\lambda$-filtered colimits and has a left adjoint $(-)\cdot C\colon\mathcal{V}\to\mathcal{K}$. (2) Similarly, the $\mathcal{V}$-functor $C\pitchfork(-)\colon\mathcal{K}\to\mathcal{K}$ preserves $\lambda$-filtered colimits (since $C$ is $\lambda$-presentable) and has a left adjoint $C\cdot(-)\colon\mathcal{K}\to\mathcal{K}$. Thus purity is preserved thanks to the proposition above. ∎
In the proposition below we denote by $\mathcal{E}_{\lambda}$ the set of arrows in $\mathcal{E}$ whose domain and codomain are $\lambda$-presentable in $\mathcal{V}$, and by $\mathcal{E}_{\lambda}^{\perp}$ the class of morphisms which are right orthogonal to every morphism in $\mathcal{E}_{\lambda}$. Then we have that $\mathcal{M}$ is closed in $\mathcal{V}^{\to}$ under $\lambda$-filtered colimits if and only if $\mathcal{E}_{\lambda}^{\perp}=\mathcal{M}$.
@@
Since $\mathcal{M}$ is closed under limits in $\mathcal{V}^{\to}$, it is enough to prove the result for $C\in\mathcal{K}_{\lambda}$. Furthermore, by Corollary 3.11 above, it is enough to show that any $(\lambda,\mathcal{E})$-pure morphism $f\colon K\to L$ in $\mathcal{V}$ lies in $\mathcal{M}$. By the hypotheses on $\mathcal{M}$, to prove that $f\in\mathcal{M}$ it is enough to show that it is right orthogonal to every morphism in $\mathcal{E}_{\lambda}$; for that it suffices to prove that the square below is a pullback.
@@
The equality $\mathcal{E}_{\lambda}^{\perp}=\mathcal{M}$ holds whenever the closure of $\mathcal{E}_{\lambda}$ in $\mathcal{E}$ under $\lambda$-filtered colimits is $\mathcal{E}$ itself. That is the case for example when $\mathcal{M}$ is closed under $\lambda$-filtered colimits in $\mathcal{V}^{\to}$ and $\mathcal{V}_{\lambda}$ is closed under $(\mathcal{E},\mathcal{M})$-factorizations in $\mathcal{V}$.
@@
    If $\mathcal{V}$ is a regular locally $\lambda$-presentable category and $\mathcal{V}_{\lambda}$ is regularly embedded in $\mathcal{V}$, the assumptions of 3.12 hold for the (regular epi, mono) factorization system since the class of monomorphisms are closed under $\lambda$-filtered colimits in $\mathcal{V}^{\to}$ and $[e,X]$ is a monomorphisms whenever $e$ is a (regular) epimorphism.
@@
    The assumptions of 3.12 are also valid in $\operatorname{\bf CMet}$ for $\mathcal{E}=$ dense. The fact that $(\omega,\mathcal{E})$-pure morphisms are isometries, which follows from 3.12, was proved in [24, 4.11].
@@
    Proposition 3.12 above does not hold in general for any factorization system. If we take $\mathcal{E}=\mathcal{K}^{\to}$ and $\mathcal{M}=\mathcal{K}^{\cong}$(see 3.4(2), then any morphism is $(\lambda,\mathcal{E})$-pure but $\mathcal{K}(C,f)$ is an isomorphism for any $C\in\mathcal{K}$ if and only if $f$ is. This factorization system satisfies $\mathcal{E}_{\lambda}^{\perp}=\mathcal{M}$, but not the requirement that $[e,X]\in\mathcal{M}$ for any $e\in\mathcal{E}$ and $X\in\mathcal{V}$.
Given a class of weights $\Psi$, following [15], we say that $\Psi$-limits are $\mathcal{E}$-stable if $\mathcal{E}$ is closed under them in $\mathcal{V}^{\to}$. In the result below the $\mathcal{E}$-stable limits can be considered to be weighted if $(\mathcal{E},\mathcal{M})$ is an enriched factorization system, and are conical otherwise.
@@
The class of $(\lambda,\mathcal{E})$-pure morphisms in a $\mathcal{V}$-category $\mathcal{K}$ is closed under $\mathcal{E}$-stable limits.
@@
Given $g\colon A\to B$, in the setting of Notation 3.1, the construction of $\mathcal{P}(g,L)$ defines a functor $\mathcal{P}(g,-)\colon\mathcal{K}_{0}\to\mathcal{V}$ which preserves $\mathcal{E}$-stable limits; this is the case since the orthogonal factorization is functorial and both $\mathcal{M}$ and $\mathcal{E}$ are closed under $\mathcal{E}$-stable limits in $\mathcal{V}^{\to}$. The same holds when considering $\mathcal{P}(g,f)$: we obtain a functor $\mathcal{P}(g,-)^{\prime}\colon\mathcal{K}_{0}^{\to}\to\mathcal{V}$ which again preserves $\mathcal{E}$-stable limits (since it is obtained by pulling back along $\mathcal{P}(g,-)$). It follows that if $f\colon K\to L$ is an $\mathcal{E}$-stable limit of some $f_{i}$, then also $r_{f}$ is the $\mathcal{E}$-stable limit of the $r_{f_{i}}$; and if each of the $r_{f_{i}}$ is in $\mathcal{E}$, then also $r_{f}$ is. Thus $\mathcal{E}$-stable limits of $(\lambda,\mathcal{E})$-pure morphisms are $(\lambda,\mathcal{E})$-pure. ∎
Recall that, given a collection $A_{i}$, $i\in I$ of objects of a locally presentable category $\mathcal{K}$ and an ultrafilter $\mathcal{U}$ on the set $I$, then the ultraproduct $\prod_{\mathcal{U}}A_{i}$ is the colimit of the directed diagram defined by the products $\prod_{i\in U}A_{i}$, where $U$ ranges through $\mathcal{U}$, and the projections $\prod_{i\in U}A_{i}\to\prod_{i\in U^{\prime}}A_{i}$, where $U^{\prime}\subseteq U$ (see, e.g., [2]).
@@
$(1)$. The ultraproduct of a family of morphisms $(f_{i})_{i}$ in $\mathcal{K}$ can be computed as a filtered colimit of products of subfamilies of the $(f_{i})_{i}$. Thus it is enough to apply the proposition above and Proposition 3.8. $(2)$. The inclusion $d_{K}\colon K\to K^{\mathcal{U}}$ is a filtered colimit of split monomorphisms, thus it is pure in the ordinary sense. By Corollary 3.9 the inclusion is also $(\omega,\mathcal{E})$-pure. $(3)$. If $f\colon K\to L$ is $(\omega,\mathcal{E})$-pure then so is $f^{\mathcal{U}}$ by point $(1)$. Conversely, if $f^{\mathcal{U}}$ is $(\omega,\mathcal{E})$-pure then also the composite $f^{\mathcal{U}}d_{K}$ is by point $(2)$ and Proposition 3.6. But $f^{\mathcal{U}}d_{K}=d_{L}f$; therefore $f$ is $(\omega,\mathcal{E})$-pure by Proposition 3.7. ∎
@@
lies in $\mathcal{E}$. Given a set of morphisms $\mathcal{H}$ in $\mathcal{K}$, we denote by $\textnormal{Inj}_{\mathcal{E}}\mathcal{H}$ the full subcategory of $\mathcal{K}$ spanned by the object which are $f$-injective over $\mathcal{E}$ for any $f\in\mathcal{H}$. We call $\mathcal{E}$-injectivity class any full subcategory of $\mathcal{K}$ arising that way; if the morphisms in $\mathcal{H}$ have $\lambda$-presentable domain and codomain, we call $\textnormal{Inj}_{\mathcal{E}}\mathcal{H}$ a $(\lambda,\mathcal{E})$-injectivity class. Then we can prove the following:
@@
    $(\lambda,\mathcal{E})$-pure subobjects.
@@
(1) is given by [15, Proposition 2.6] and (2) by [15, Proposition 2.7] since $\mathcal{E}$, being part of a factorization system, is closed in $\mathcal{V}^{\to}$ under all conical colimits. (3) Take a $(\lambda,\mathcal{E})$-pure morphism $f\colon K\to L$ with $L\in\textnormal{Inj}_{\mathcal{E}}\mathcal{H}$. Let $g\colon A\to B$ be in $\mathcal{H}$, which is then between $\lambda$-presentable objects. By $\mathcal{E}$-injectivity of $L$ with respect to $g$, the map $\mathcal{K}(g,L)$ is in $\mathcal{E}$, so that $\mathcal{P}(g,L)\cong\mathcal{K}(A,L)$. It follows that $\mathcal{P}(g,f)\cong\mathcal{K}(A,K)$, it being the pullback of $\mathcal{P}(g,L)$ along $\mathcal{K}(A,f)$. Therefore the map $r$ appearing in the definition of $(\lambda,\mathcal{E})$-purity coincides with $\mathcal{K}(g,K)\colon\mathcal{K}(B,K)\to\mathcal{K}(A,K)$. Since $f$ is $(\lambda,\mathcal{E})$-pure, $\mathcal{K}(g,K)$ is in $\mathcal{E}$ and hence $K$ is $\mathcal{E}$-injective with respect to $g$. ∎
@@
Suppose that $(\mathcal{E},\mathcal{M})$ is proper, $\mathcal{M}$ is closed under $\lambda$-filtered colimits in $\mathcal{V}^{\to}$, and that products are $\mathcal{E}$-stable. Then every $(\lambda,\mathcal{E})$-injectivity class $\mathcal{L}$ in a locally $\lambda$-presentable $\mathcal{V}$-category $\mathcal{K}$ is an ordinary $\lambda$-injectivity class. Indeed, by the proposition above $\mathcal{L}$ is closed under products, $\lambda$-filtered colimits and $\lambda$-pure subobjects (see 8.12). Following [27], it is then a $\lambda$-injectivity class.
@@
Let $\mathcal{W}=(\mathcal{W}_{0},I_{\mathcal{W}},\otimes_{\mathcal{W}})$ be locally $\lambda$-presentable as closed category, and $\mathcal{V}$ be a locally $\lambda$-presentable $\mathcal{W}$-category. Assume moreover that $\mathcal{V}_{0}$ is endowed with an enriched symmetric monoidal closed structure $(\mathcal{V},I_{\mathcal{V}},\otimes_{\mathcal{V}})$ which makes it locally $\lambda$-presentable as a closed category. Then the $\mathcal{W}$-functor $U:=\mathcal{V}(I_{\mathcal{V}},-)\colon\mathcal{V}\to\mathcal{W}$ is continuous and preserves $\lambda$-filtered colimits; by local presentability of $\mathcal{V}$ and $\mathcal{W}$, it has a left adjoint $F\colon\mathcal{W}\to\mathcal{V}$ which is strong monoidal ([11, 2.1]). Under these assumptions we have a change of base functor
@@
which sends a $\mathcal{V}$-category $\mathcal{K}$ to the $\mathcal{W}$-category $U_{*}\mathcal{K}$ with the same objects as $\mathcal{K}$ and homs $(U_{*}\mathcal{K})(X,Y):=U\mathcal{K}(X,Y)$, so that $\mathcal{K}_{0}=(U_{*}\mathcal{K})_{0}$ and hence morphisms in $\mathcal{K}$ are the same as morphisms in $U_{*}\mathcal{K}$. See for instance [7, Section 6]. It is now easy to see, directly from the definition of $(\lambda,\mathcal{E})$-purity (since $U$ preserves pullbacks) that the following properties hold whenever $\mathcal{V}$ and $\mathcal{W}$ are equipped with factorization systems $(\mathcal{E}_{\mathcal{V}},\mathcal{M}_{\mathcal{V}})$ and $(\mathcal{E}_{\mathcal{W}},\mathcal{M}_{\mathcal{W}})$ respectively:
@@
A direct application of the result above is obtained by considering $\operatorname{\bf Ban}$ as a $\operatorname{\bf CMet}$-category (see [3, Example 4.5(3)]); the induced $\operatorname{\bf CMet}$-functor $U$ is the unit-ball functor $U\colon\operatorname{\bf Ban}\to\operatorname{\bf Met}$. Moreover, the (dense, isometry) factorization system on $\operatorname{\bf Ban}$ can be obtained, as in (3) above, by taking the preimage of the factorization system (dense, isometry) in $\operatorname{\bf CMet}$. It follows that a morphism $f$ in a $\operatorname{\bf Ban}$-category $\mathcal{K}$ is $(\lambda,\textnormal{dense})$-pure if and only if it is so in the $\operatorname{\bf CMet}$-category $U_{*}\mathcal{K}$.
@@
Following the nomenclature of [18] for accessible $\mathcal{V}$-categories, we know that an accessibly embedded full subcategory of a locally presentable $\mathcal{V}$-category is accessible if and only if it is closed under ordinary $\lambda$-pure subobjects for some $\lambda$ (see [18, 3.23] and [2, 2.36]). The same holds with respect to conical accessibility. Let us consider now enriched purity. Under the hypothesis of Corollary 3.9 and using Proposition 3.17 above, any accessibly embedded full subcategory of a locally presentable $\mathcal{V}$-category that is closed under $(\lambda,\mathcal{E})$-pure subobjects, for some $\lambda$, is accessible. We do not know, however, whether the converse implication holds; that is, if every accessible and accessibly embedded $\mathcal{V}$-category in a locally presentable $\mathcal{V}$-category is closed under enriched pure subobjects.

### v3

@@ 3. Enriched purity
Let $f\colon K\to L$ be a $\lambda$-filtered colimit in $\mathcal{K}^{\to}$ of some $(\lambda,\mathcal{E})$-pure morphisms $f_{i}\colon K_{i}\to L_{i}$, and consider any $g\colon A\to B$ in $\mathcal{K}_{\lambda}$. Since $\mathcal{K}(A,K)$ is the $\lambda$-filtered colimit of $\mathcal{K}(A,K_{i})$ and $\mathcal{K}(B,K)$ is the $\lambda$-filtered colimit of $\mathcal{K}(B,K_{i})$, the map
$\mathcal{K}(g,K)\colon\mathcal{K}(B,K)\to\mathcal{K}(A,K)$
is the $\lambda$-filtered colimit of the maps
$\mathcal{K}(g,K_{i})\colon\mathcal{K}(B,K_{i})\to\mathcal{K}(A,K_{i});$
the same applies for $L$ in place of $K$. Now, by $(\lambda,\mathcal{E})$-purity of each $f_{i}$, the $(\mathcal{E},\mathcal{M})$-factorization of each $\mathcal{K}(g,K_{i})$ is given by the object $\mathcal{P}(g,f_{i})$, see Remark 3.3. By orthogonality of the factorization system, we obtain a $\lambda$-filtered diagram on the $\mathcal{P}(g,f_{i})$ whose colimit $X(g,f)\in\mathcal{V}$ induces the $(\mathcal{E},\mathcal{M})$-factorization of $\mathcal{K}(g,K)$; this is because both $\mathcal{E}$ and $\mathcal{M}$ are closed under $\lambda$-filtered colimits in $\mathcal{K}^{\to}$. But $\lambda$-filtered colimits commute with pullbacks in $\mathcal{V}$; thus $X(g,f)$ is also the pullback along $\mathcal{K}(A,f)$ of the colimit of $\mathcal{P}(g,L_{i})$ Since this colimit coincides with $\mathcal{P}(g,L)$ then $X(g,f)\cong\mathcal{P}(g,f)$ and thus $f$ is $(\lambda,\mathcal{E})$-pure with respect to $g$. ∎
@@
Let $L\colon\mathcal{L}\to\mathcal{K}$ be the left adjoint to $F$; since $F$ preserves $\lambda$-filtered colimits then $L$ preserves the $\lambda$-presentable objects. Consider now a $(\lambda,\mathcal{E})$-pure morphism $f\colon K\to L$ in $\mathcal{K}$ and any morphism $g\colon A\to B$ in $\mathcal{L}$ between $\lambda$-presentable objects. Then $\mathcal{L}(g,FL)\cong\mathcal{K}(Lg,L)$ and $\mathcal{L}(A,Ff)\cong\mathcal{K}(LA,f)$, so that
@@
as well as $r_{g,Ff}\cong r_{Lg,f}$ in $\mathcal{V}^{\to}$, where indices denote to which pair of morphisms Notation 3.1 applies. Now, since $A$ and $B$ are $\lambda$-presentable in $\mathcal{L}$, the morphism $Lg$ has $\lambda$-presentable domain and codomain in $\mathcal{K}$; thus we can apply the $(\lambda,\mathcal{E})$-purity of $f$ in $\mathcal{K}$ to deduce that $r_{Lg,f}$ lies in $\mathcal{E}$. Since that is isomorphic to $r_{g,Ff}$, the $(\lambda,\mathcal{E})$-purity of $Ff$ follows. ∎
@@
(1) It follows by Proposition 3.11 above since $\mathcal{K}(C,-)$ preserves $\lambda$-filtered colimits and has a left adjoint $(-)\cdot C\colon\mathcal{V}\to\mathcal{K}$.
(2) Similarly, the $\mathcal{V}$-functor $C\pitchfork(-)\colon\mathcal{K}\to\mathcal{K}$ preserves $\lambda$-filtered colimits (since $C$ is $\lambda$-presentable) and has a left adjoint $C\cdot(-)\colon\mathcal{K}\to\mathcal{K}$. Thus purity is preserved thanks to Proposition 3.11 above. ∎
In the proposition below we denote by $\mathcal{E}_{\lambda}$ the set of arrows in $\mathcal{E}$ whose domain and codomain are $\lambda$-presentable in $\mathcal{V}$, and by $\mathcal{E}_{\lambda}^{\perp}$ the class of morphisms which are right orthogonal to every morphism in $\mathcal{E}_{\lambda}$.
@@
Since $\mathcal{M}$ is closed under limits in $\mathcal{V}^{\to}$, it is enough to prove the result for $C\in\mathcal{K}_{\lambda}$. Furthermore, by Corollary 3.12 above, it is enough to show that any $(\lambda,\mathcal{E})$-pure morphism $f\colon K\to L$ in $\mathcal{V}$ lies in $\mathcal{M}$.
By the hypotheses on $\mathcal{M}$, to prove that $f\in\mathcal{M}$ it is enough to show that it is right orthogonal to every morphism in $\mathcal{E}_{\lambda}$; for that it suffices to prove that the square below is a pullback.
@@
The equality $\mathcal{E}_{\lambda}^{\perp}=\mathcal{M}$ holds whenever the closure of $\mathcal{E}_{\lambda}$ in $\mathcal{E}$ under $\lambda$-filtered colimits is $\mathcal{E}$ itself (because $\mathcal{E}$ is closed under colimits). That is the case for example when $\mathcal{M}$ is closed under $\lambda$-filtered colimits in $\mathcal{V}^{\to}$ and $\mathcal{V}_{\lambda}$ is closed under $(\mathcal{E},\mathcal{M})$-factorizations in $\mathcal{V}$.
@@
    If $\mathcal{V}$ is a regular locally $\lambda$-presentable category and $\mathcal{V}_{\lambda}$ is regularly embedded in $\mathcal{V}$, the assumptions of 3.13 hold for the (regular epi, mono) factorization system since the class of monomorphisms are closed under $\lambda$-filtered colimits in $\mathcal{V}^{\to}$ and $[e,X]$ is a monomorphisms whenever $e$ is a (regular) epimorphism.
@@
    The assumptions of 3.13 are also valid in $\operatorname{\bf CMet}$ for $\mathcal{E}=$ dense. The fact that $(\omega,\mathcal{E})$-pure morphisms are isometries, which follows from 3.13, was proved in [24, 4.11].
@@
    Proposition 3.13 above does not hold in general for any factorization system. If we take $\mathcal{E}=\mathcal{K}^{\to}$ and $\mathcal{M}=\mathcal{K}^{\cong}$(see 3.5(2), then any morphism is $(\lambda,\mathcal{E})$-pure but $\mathcal{K}(C,f)$ is an isomorphism for any $C\in\mathcal{K}$ if and only if $f$ is. This factorization system satisfies $\mathcal{E}_{\lambda}^{\perp}=\mathcal{M}$, but not the requirement that $[e,X]\in\mathcal{M}$ for any $e\in\mathcal{E}$ and $X\in\mathcal{V}$.
Below, following [15], we say that a limit of some shape is $\mathcal{E}$-stable if $\mathcal{E}$ is closed under such limits in $\mathcal{V}^{\to}$.
@@
The class of $(\lambda,\mathcal{E})$-pure morphisms in a $\mathcal{V}$-category $\mathcal{K}$ is stable under $\mathcal{E}$-stable (conical) limits in $\mathcal{K}^{\to}$.
@@
Assume that $f\colon K\to L$ is the $\mathcal{E}$-stable limit of some $(\lambda,\mathcal{E})$-pure morphisms $f_{i}\colon K_{i}\to L_{i}$ in $\mathcal{K}^{\to}$. For any $g\colon A\to B$ between $\lambda$-presentable objects, we can consider the diagram of Notation 3.1 applied to $f_{i}$ and $g$:
$\mathcal{K}(B,K_{i})$$\mathcal{K}(B,L_{i})$$\mathcal{K}(A,K_{i})$$\mathcal{K}(A,L_{i})$$\mathcal{P}(g,f_{i})$$\mathcal{P}(g,L_{i})$$\mathcal{K}(B,f_{i})$$r^{\prime}_{i}$$\mathcal{K}(A,f_{i})$$\mathcal{K}(g,K_{i})$$r_{i}$
Since the factorization system $(\mathcal{E},\mathcal{M})$ is functorial, $\mathcal{E}$ and $\mathcal{M}$ are stable under $\mathcal{E}$-stable limits, pullbacks commute with all limits, and the representables preserve all limits, it follows that the componentwise limit of the diagram above in $\mathcal{V}$, is exactly the diagram displayed in Notation 3.1 for $f$ and $g$. It follows in particular that
$r\colon\mathcal{K}(B,K)\to\mathcal{P}(g,f)$
is the $\mathcal{E}$-stable limit of the $r_{i}$’s. But each $r_{i}$ lies in $\mathcal{E}$, by purity of $f_{i}$; thus also $r$ lies in $\mathcal{E}$ and $f$ is $(\lambda,\mathcal{E})$-pure. ∎
Recall that, given a collection $A_{i}$, $i\in I$ of objects of a locally presentable category $\mathcal{K}$ and an ultrafilter $\mathcal{U}$ on the set $I$, then the ultraproduct $\prod_{\mathcal{U}}A_{i}$ is defined as the colimit of the directed diagram induced by the products $\prod_{i\in U}A_{i}$, where $U$ ranges through $\mathcal{U}$, and the projections $\prod_{i\in U}A_{i}\to\prod_{i\in U^{\prime}}A_{i}$, where $U^{\prime}\subseteq U$ (see, e.g., [2]).
@@
$(1)$. The ultraproduct of a family of morphisms $(f_{i})_{i}$ in $\mathcal{K}$ can be computed as a filtered colimit of products of subfamilies of the $(f_{i})_{i}$. Thus it is enough to apply Proposition 3.16 above and Proposition 3.9.
$(2)$. The inclusion $d_{K}\colon K\to K^{\mathcal{U}}$ is a filtered colimit of split monomorphisms, thus it is pure in the ordinary sense. By Corollary 3.10 the inclusion is also $(\omega,\mathcal{E})$-pure.
$(3)$. If $f\colon K\to L$ is $(\omega,\mathcal{E})$-pure then so is $f^{\mathcal{U}}$ by point $(1)$. Conversely, if $f^{\mathcal{U}}$ is $(\omega,\mathcal{E})$-pure then also the composite $f^{\mathcal{U}}d_{K}$ is by point $(2)$ and Proposition 3.7. But $f^{\mathcal{U}}d_{K}=d_{L}f$; therefore $f$ is $(\omega,\mathcal{E})$-pure by Proposition 3.8. ∎
@@
lies in $\mathcal{E}$. Given a set of morphisms $\mathcal{H}$ in $\mathcal{K}$, we denote by $\textnormal{Inj}_{\mathcal{E}}\mathcal{H}$ the full subcategory of $\mathcal{K}$ spanned by the object which are $f$-injective over $\mathcal{E}$ for any $f\in\mathcal{H}$. We call $\mathcal{E}$-injectivity class any full subcategory of $\mathcal{K}$ arising that way; if the morphisms in $\mathcal{H}$ have $\lambda$-presentable domain and codomain, we call $\textnormal{Inj}_{\mathcal{E}}\mathcal{H}$ a $(\lambda,\mathcal{E})$-injectivity class.
Then we can prove the following result. Note that below we use the term “subobject” when referring to the closure property under $(\lambda,\mathcal{E})$-pure morphism even though such morphisms need not be monomorphic in general; we do this since that is the standard way of phrasing the closure property.
@@
    $(\lambda,\mathcal{E})$-pure subobjects; that is, if $f\colon K\to L$ is $(\lambda,\mathcal{E})$-pure and $L\in\textnormal{Inj}_{\mathcal{E}}\mathcal{H}$, then also $K\in\textnormal{Inj}_{\mathcal{E}}\mathcal{H}$.
@@
(1) is given by [15, Proposition 2.6] and (2) by [15, Proposition 2.7] since $\mathcal{E}$, being part of a factorization system, is closed in $\mathcal{V}^{\to}$ under all conical colimits.
(3) Take a $(\lambda,\mathcal{E})$-pure morphism $f\colon K\to L$ with $L\in\textnormal{Inj}_{\mathcal{E}}\mathcal{H}$. Let $g\colon A\to B$ be in $\mathcal{H}$, which is then between $\lambda$-presentable objects.
By $\mathcal{E}$-injectivity of $L$ with respect to $g$, the map $\mathcal{K}(g,L)$ is in $\mathcal{E}$, so that $\mathcal{P}(g,L)\cong\mathcal{K}(A,L)$. It follows that $\mathcal{P}(g,f)\cong\mathcal{K}(A,K)$, it being the pullback of $\mathcal{P}(g,L)$ along $\mathcal{K}(A,f)$. Therefore the map $r$ appearing in the definition of $(\lambda,\mathcal{E})$-purity (Notation 3.1) coincides with
$\mathcal{K}(g,K)\colon\mathcal{K}(B,K)\to\mathcal{K}(A,K).$
Since $f$ is a $(\lambda,\mathcal{E})$-pure morphism, $\mathcal{K}(g,K)$ lies in $\mathcal{E}$ and hence $K$ is $\mathcal{E}$-injective with respect to $g$. ∎
@@
Suppose that $(\mathcal{E},\mathcal{M})$ is proper, $\mathcal{M}$ is closed under $\lambda$-filtered colimits in $\mathcal{V}^{\to}$, and that products are $\mathcal{E}$-stable. Then every $(\lambda,\mathcal{E})$-injectivity class $\mathcal{L}$ in a locally $\lambda$-presentable $\mathcal{V}$-category $\mathcal{K}$ is an ordinary $\lambda$-injectivity class.
Indeed, by the proposition above $\mathcal{L}$ is closed under products, $\lambda$-filtered colimits and $\lambda$-pure subobjects (see 8.12). Following [27], it is then a $\lambda$-injectivity class.
@@
Let $\mathcal{W}=(\mathcal{W}_{0},I_{\mathcal{W}},\otimes_{\mathcal{W}})$ be locally $\lambda$-presentable as closed category, and $\mathcal{V}$ be a locally $\lambda$-presentable $\mathcal{W}$-category. Assume moreover that $\mathcal{V}_{0}$ is endowed with an enriched symmetric monoidal closed structure $(\mathcal{V},I_{\mathcal{V}},\otimes_{\mathcal{V}})$ which makes it locally $\lambda$-presentable as a closed category. Then the $\mathcal{W}$-functor
$U:=\mathcal{V}(I_{\mathcal{V}},-)\colon\mathcal{V}\to\mathcal{W}$
is continuous and preserves $\lambda$-filtered colimits; by local presentability of $\mathcal{V}$ and $\mathcal{W}$, it has a left adjoint $F\colon\mathcal{W}\to\mathcal{V}$ which is strong monoidal ([11, 2.1]).
Under these assumptions there is a change of base functor
@@
which sends a $\mathcal{V}$-category $\mathcal{K}$ to the $\mathcal{W}$-category $U_{*}\mathcal{K}$ with the same objects as $\mathcal{K}$ and homs
$(U_{*}\mathcal{K})(X,Y):=U\mathcal{K}(X,Y),$
so that $\mathcal{K}_{0}=(U_{*}\mathcal{K})_{0}$ and hence morphisms in $\mathcal{K}$ are the same as morphisms in $U_{*}\mathcal{K}$. See for instance [8, Section 6].
It is now easy to see that the following properties hold whenever $\mathcal{V}$ and $\mathcal{W}$ are equipped with factorization systems $(\mathcal{E}_{\mathcal{V}},\mathcal{M}_{\mathcal{V}})$ and $(\mathcal{E}_{\mathcal{W}},\mathcal{M}_{\mathcal{W}})$ respectively:
@@
This follows directly from the definition of $(\lambda,\mathcal{E})$-purity (since $U$ preserves pullbacks) by applying $U$ to the diagram of Notation 3.1. We leave the details to the reader as we will not makes use this result in any of our Theorems.
@@
A direct application of the result above is obtained by considering $\operatorname{\bf Ban}$ as a $\operatorname{\bf CMet}$-category (see [3, Example 4.5(3)]); the induced $\operatorname{\bf CMet}$-functor $U$ is the unit-ball functor $U\colon\operatorname{\bf Ban}\to\operatorname{\bf Met}$. Moreover, the (dense, isometry) factorization system on $\operatorname{\bf Ban}$ can be obtained, as in (3) above, by taking the preimage of the factorization system (dense, isometry) in $\operatorname{\bf CMet}$.
It follows that a morphism $f$ in a $\operatorname{\bf Ban}$-category $\mathcal{K}$ is $(\lambda,\textnormal{dense})$-pure if and only if it is so in the $\operatorname{\bf CMet}$-category $U_{*}\mathcal{K}$.
@@
Following the nomenclature of [18] for accessible $\mathcal{V}$-categories, we know that an accessibly embedded full subcategory of a locally presentable $\mathcal{V}$-category is accessible if and only if it is closed under ordinary $\lambda$-pure subobjects for some $\lambda$ (see [18, 3.23] and [2, 2.36]). The same holds with respect to conical accessibility.
Let us consider now enriched purity. Under the hypothesis of Corollary 3.10 and using Proposition 3.18 above, any accessibly embedded full subcategory of a locally presentable $\mathcal{V}$-category that is closed under $(\lambda,\mathcal{E})$-pure subobjects, for some $\lambda$, is accessible. We do not know, however, whether the converse implication holds; that is, if every accessible and accessibly embedded $\mathcal{V}$-category in a locally presentable $\mathcal{V}$-category is closed under enriched pure subobjects.